A mixture of krypton and xenon gases at a total pressure of 874 mm Hg contains krypton at a partial pressure of 508 mm Hg. If the gas mixture contains 14.6 grams of krypton, how many grams of xenon are present?



1
Expert's answer
2021-05-03T07:24:31-0400

Krypton +Xenon= Ptotal

Ptotal=874mmHg

Moles of Krypton gas=14.6g×

"\frac{1mol}{83.8g/mol}"

="0.174moles"

Partial pressure of Krypton=Mole fraction ×total pressure

Mole fraction of Krypton="\frac{874mmHg}{508mmHg}"


"=1.720moles"

But total moles=moles×mole fraction

=0.174×1.720="0.3moles"

Total moles=0.3moles

But Partial pressure of Xenon=(874mmHg-508mmHg)=366mmHg

Mole fraction of Xenon gas=

"\frac{874mmHg}{366mmHg}"

=2.388moles

Moles of Xenon gas =(2.388×0.3)=0.716moles

But Moles of Xenon gas=mass ×"\frac{1mole Xe}{131.29g/mole}"

Moles of Xenon gas=0.716moles

Mass of Xenon gas="\frac{0.716}{0.01}"


="71.6g"